OpenStudy (anonymous):

A jar contains nine blue marbles, eight yellow marbles, 11 black marbles, and 12 red marbles. Without looking, what is the probability of pulling a yellow marble from the jar? 1/5 8/40 3/10 9/40

OpenStudy (anonymous):

8/40

OpenStudy (anonymous):

8 yellow marbles 40 total marbles

OpenStudy (marissalovescats):

Well there are 9 blue marbles, 8 yellow, 11 black, and 12 red. So 9+8+11+12=40 so there are 40 total marbles. And the probability is pulling 8 yellow marbles out of 40 because there are 8. Which simplifies to 1/5

OpenStudy (anonymous):

yup if supposed to simplify 8/40 = 1/5